OpenStudy (anonymous):

I am confused help! What linear systems of equation does this matrix represent? [2 -3|10] [5 6|-12] Except there are two big "[" on the outside of them.

OpenStudy (anonymous):

2 * x1 - 3 *x2 = 10 and 5 * x1 + 6*x2 = -12

ganeshie8 (ganeshie8):

\[ \left[ \begin{array}{ccc} 2 & -3 & 10 \\ 5 & 6 & -12 \end{array} \right]\]

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

YEah

OpenStudy (anonymous):

I am bad at math.........

ganeshie8 (ganeshie8):

if we say x and y are the variables the, the equation wiud be :- 2x - 3y = 10 5x + 6y = -12

ganeshie8 (ganeshie8):

each row in that matrix represents an equation, thats all to it

OpenStudy (anonymous):

Ohh it's that easy? I was over thinking this

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

Thank you

ganeshie8 (ganeshie8):

np :)